OpenStudy (anonymous):

ln 1/sqrt e = -1/2 write in exponential form pls show steps ln 1/sqrt e = -1/2 write in exponential form pls show steps @Mathematics

OpenStudy (anonymous):

is e just a variable? or is it the constant e? because i see this: \[\frac{1}{\sqrt{e}} = \frac{-1}{2}\]